# 2Meituan, Shanghai, China

### v1

@@ Abstract.
Recent advances in code agents have enabled automated software development at the project level, supported by large language models (LLMs) and widely adopted tools. However, existing benchmarks for code agent evaluation face two major limitations: high annotation cost and expertise requirements, and rigid evaluation metrics that rely primarily on unit tests. To address these challenges, we propose an agent-driven benchmark construction pipeline that leverages human supervision to efficiently generate diverse and challenging project-level tasks. Based on this approach, we introduce PRDBench, a novel benchmark comprising 50 real-world Python projects across 20 domains, each with structured Product Requirement Document (PRD) requirements, comprehensive evaluation criteria, and reference implementations. PRDBench features rich data sources, high task complexity, and flexible metrics. We further employ an Agent-as-a-Judge paradigm to score agent outputs, enabling the evaluation of various test types beyond unit tests. Extensive experiments on PRDBench demonstrate its effectiveness in assessing the capabilities of both code agents and evaluation agents, providing a scalable and robust framework for annotation and evaluation.
@@
Code Agent, Agent Evaluation, Large Language Models
* corresponding author.
{CCSXML}
[…]
@@ 1. Introduction
In recent years, code agents have made significant progress, capable of solving increasingly complex programming tasks. From initially focusing on single-file code generation to now advancing towards complete project-level software development, Code Agents are demonstrating powerful software development capabilities. LLM-based Code Agents such as CursorAgent Inc. (2025), Claude Code Anthropic (2025a) and Gemini CLI Mullen and Salva (2025) have been widely adopted in real-world development scenarios, accelerating the proliferation of automated programming tools.
As code agents continue to advance, a variety of benchmarks have emerged to evaluate their capabilities. Several works Chan et al. (2024); Huang et al. (2024); Daghighfarsoodeh et al. (2025) focus on assessing LLMs and code agents in data science tasks, including subtasks such as data processing and machine learning training. Other benchmarks Jimenez et al. (2023); Fu et al. (2025); Vergopoulos et al. (2025) select complex code repositories from GitHub pull requests and transform them, often with minimal human intervention, into testable scenarios. In addition, PaperBench Starace et al. (2025) evaluates the ability of agents to reproduce paper-level code implementations, providing a more rigorous assessment of their research reproducibility.
Despite these advancements, current benchmarks face two major limitations. First, the creation of high-quality evaluation datasets requires substantial human effort and expertise. Generating reliable unit tests or grading tasks often depends on expert annotators, especially when migrating real-world scenarios such as Kaggle competitions or GitHub pull requests. These tasks demand domain-level verification and annotation to produce usable test cases. As code agents advance rapidly, the requirements for annotators’ expertise and time investment continue to increase. For example, PaperBench Starace et al. (2025) recruited ICML authors as annotators, with each annotation task demanding several days of work from PhD-level experts. This process not only incurs high annotation costs, but also restricts the diversity of evaluation data, as recruiting experts from various domains is challenging and often leads to datasets being sourced from a single domain.
Second, the rigidity of the evaluation limits the broad applicability of current benchmarks. Existing benchmarks Jimenez et al. (2023) primarily rely on unit test pass rates to assess code agent performance. While unit tests are effective for verifying specific functions or components, they are limited in scope and cannot fully cover the diverse testing needs encountered in project-level software development. In practical scenarios, quality assurance involves not only unit testing, but also integration testing, end-to-end testing, performance testing, and security testing, which are all essential for robust software delivery. Relying solely on unit tests is not only insufficient, but also overly restrictive, as it enforces strict requirements on project interfaces and implementation details. Many real-world tasks require broader validation methods. To address these gaps, it is necessary to introduce additional evaluation strategies such as shell interaction and file comparison, enabling more comprehensive and flexible assessment of code agents in complex development environments.
@@
SWE-Bench Jimenez et al. (2023) Pull Requests 12 Web Crawler 2,294 Unit Test 70.3%
MLEBench Chan et al. (2024) AI Competition 75 Human 75 Test Set 51.1%
DevAI Zhuge et al. (2024) AI Developement 55 Human 365 Agent 73.0%
PaperBench Starace et al. (2025) AI Research 20 Human (PhD.) 8,316 Human/LLM 21.0%
PRDBench Engineering Development 50 Agent & Human 1,262 Agent 45.5%
Table 1. Comparison between PRDBench and project-level code agent benchmarks. We employ the performance of code agents with Claude as a benchmark indicator for assessing the task difficulty of our proposed evaluation suite. The results from external report Anthropic (2025a) is denoted by *, from our implementation is denoted by $\dagger$, otherwise we use the best results from the original paper with author specified Claude-driven agents.
@@
Based on this approach, we build a Product Requirement Document (PRD)-centered benchmark, named PRDBench [^1] Code of PRDBench Evaluation is available in https://github.com/AGI-Eval-Official/PRDBench. PRDBench data will be available soon.. PRDBench consists of 50 real-world coding tasks, each defined by a structured PRD, an well-specified and verifiable criteria scheme, and a standard solution code repository. The coding tasks are sourced from real-world requirements, academic projects, and thesis work, spanning 20 common subdomains. For each task, the PRD provides a criteria scheme that facilitates comprehensive human-like quality assurance (QA) checks. We present a comparison between PRDBench and previous code agent benchmarks in Table 1, where PRDBench offers a comprehensive, multifaceted, project-level benchmark for code agents. During evaluation, we employ the Agent-as-a-Judge paradigm to score code according to the criteria scheme, enabling the evaluation of a wide range of test types beyond unit tests.
@@
  • We adopt the Agent-as-a-Judge paradigm for evaluation, enabling flexible adaptation to human-like QA assessment beyond traditional unit tests.
  • We conduct extensive experiments on PRDBench, providing useful insights into both code agents and the evaluation agent.
@@ 2.1. Code Agent Evaluation
To keep pace with the rapid development of LLMs’ and agents’ coding capabilities, recent benchmarks increasingly emphasize the construction of executable and complex tasks within real software projects. These benchmarks are typically built by mining human-generated data from online platforms or through extensive manual annotation. Some benchmarks Li et al. (2024b); Huang et al. (2024) require annotators to manually create test points for each step, while others Jimenez et al. (2024); Yang et al. (2024); Pan et al. (2024); Wu et al. (2024); Hai et al. (2025) rely on authentic submission records, using the inherent difficulty of submitted code as a basis for evaluation. Additionally, some benchmarks Li et al. (2024a); Zhuo et al. (2025); Peng et al. (2024) leverage existing resources and require expert-driven rewriting, which demands significant manual effort. As code agents become more capable, these annotation paradigms require increasingly specialized expertise, making the recruitment of expert annotators for each domain costly and limiting the scalability and diversity of benchmark construction.
In terms of evaluation metrics, most existing benchmarks adopt relatively narrow criteria, such as unit test pass rates Jimenez et al. (2023); Li et al. (2024b); Xiang et al. (2025), which primarily assess functional correctness. ProjectEval Liu et al. (2025) simulates user interaction to evaluate projects, focusing on specific types of tasks. However, these evaluation criteria are limited to specific task types and cannot comprehensively assess diverse, complex software projects.
@@ 2.2. Agent-as-a-Judge
Due to the inherent limitations of unit test-based benchmarks, recent studies have begun to explore the use of LLMs as judges for agent evaluation, a paradigm often referred to as LLM-as-a-Judge. LLMs can provide capabilities such as image recognition Chen et al. (2024), code alignment Weyssow et al. (2024), and quesion-answering Ho et al. (2025), breaking the restriction that benchmark metrics must be executable. With the emergence of agents equipped with tools, LLMs are able to read large volumes of files and perform extensive analyses, offering new possibilities for evaluation. DevAI Zhuge et al. (2024) is the first work to apply the agent-as-a-judge paradigm to code agent evaluation. It focuses on checking program execution and output format, but lacks comprehensive assessment of overall engineering effectiveness. Some studiesWang et al. (2025), have explored LLM-based agent frameworks for evaluating code generation, improving the accuracy of Agent-as-a-Judge through architectural innovations. However, the lack of well-defined metrics and code agents has limited the use of agent-as-a-judge. PRDBench fills this gap with structured criteria and dedicated tools, enabling more robust evaluation.
@@ 3. PRDBench
In this section, we describe the construction and design of PRDBench, covering seed task filtering, human annotation, and the configuration of evaluation agents. We list education background of hired annotators in Appendix A.
@@ 3.1. Seed Tasks
Our seed tasks are sourced from a variety of real-world project requirements, including user requests from end-to-end AI product development platforms, academic theses and projects. To ensure the suitability and consistency of tasks for PRDBench, we apply a rigorous filtering process. Specifically, selected tasks must satisfy: (1) the task can be fully implemented in Python; and (2) all datasets required for the task are publicly accessible. We choose Python as the programming language for PRDBench primarily due to its versatility and comprehensive ecosystem, which supports a wide range of programming paradigms and application domains. This ensures that PRDBench tasks are representative of diverse, real-world scenarios.
@@ 3.2. Agent-Driven Data Production
After selecting seed tasks, we utilize SOTA LLMs (such as GPT-4.1) or code agents (such as Claude Code) to generate detailed and standardized PRD documents, ensuring clarity and completeness in task specification. The PRD serves as the evaluation blueprint for PRDBench, and includes sections such as Requirement Overview, Functional Requirements, and Data Requirements. Based on the PRD, we employ GPT-4.1 to generate a corresponding metric outline, structuring test cases using the Arrange-Act-Assert (AAA) methodology Wei et al. (2025). The Arrange step sets up the test case by preparing the necessary files, input data, and environment configurations. The Act step focuses on the core behavior to be tested, such as running the program or performing interaction tests to obtain output results. The Assert step verifies the expected outcomes by checking the system’s response or state, ultimately determining whether the test passes or fails. This structure is applicable to a wide range of code-related testing scenarios, providing PRDBench with a comprehensive and executable testing plan.
Step 2: Scaffold and Criteria Generation
@@
After initializing with the PRDs and metric outlines, we employ state-of-the-art code agents to generate the code scaffold for each task. The scaffold includes both module design and interface design, serving as the core framework of the entire project. In some cases, particularly for complex algorithms, the implementation may not be fully correct as long as it does not affect the annotation process. Building upon the scaffold, we further utilize code agents to expand and refine the metric outline into a specific criteria scheme, which includes necessary test interfaces and test artifacts. The presence of well-defined code interfaces in the scaffold simplifies and standardizes the generation of the criteria scheme, making the process more efficient and consistent.
Figure 2. A task example of PRDBench.
Step 3: Human Inspection
@@
With the scaffold and the criteria scheme in place, the inspection process for human annotators becomes straightforward. Annotators conduct inspection only by running the tests to verify whether the interfaces function correctly and whether the expected outputs in the criteria scheme align with the PRD requirements. Notably, even for tasks sourced from expertise projects or academic papers, annotators only need basic computer science knowledge to perform effective checking and annotation.
Step 4: Agent-Based Fix and Refinement
@@
If any issues are identified during step 3, annotators shall provide targeted feedback to the code agent for refining the scaffold or criteria scheme. The code agent then revises the relevant components, and the inspection process is repeated. This iterative process continues until all issues are resolved, ensuring the quality and correctness of the data.
Step 5: Remove Scaffold
@@
In the final stage, the scaffold is removed and only the criteria scheme, test artifacts, required data, and PRD are retained in PRDBench. The removal of scaffold ensures that the code agents to be evaluated generate the code from scratch, thereby we can assess the end-to-end development capacity of the agents.
@@
3.3. EvalAgent and Evaluation
Figure 3. Overview of PRDBench evaluation. The EvalAgent executes tests based on the criteria scheme using various tools, compares outputs (files or results) with expected outputs, and generates a report for the submitted code.
For each item in the criteria scheme, PRDBench provides executable terminal commands and expected outputs, enabling systematic verification. PRDBench employs a lightweight evaluation agent, EvalAgent, to automate the validation of code generated by code agents against predefined criteria schemes, as illustrated in Figure 3. EvalAgent is equipped with six core tools, including file reading and writing, command-line execution, image handling, and a judge tool. Among these, the dealgraph tool offers a multimodal LLM interface (GPT-4o in this paper), allowing the agent to process images by submitting both the image and a descriptive prompt to obtain results—this is particularly useful for tasks requiring visual verification. The judge tool is specifically designed for PRDBench. It accepts files containing simulated user inputs, enabling comprehensive terminal log generation for thorough analysis without the need for manual input specification. By automating the validation process, EvalAgent significantly reduces manual intervention and improves evaluation efficiency.
Within PRDBench, the EvalAgent performs three main categories of tests, which collectively simulate the different stages of QA in Python projects:

### v2

# Automatically Benchmarking LLM Code Agents through Agent-driven Annotation and Evaluation

@@ Abstract.
Recent advances in code agents have enabled automated software development at the project level, supported by large language models (LLMs). However, existing benchmarks for code agent evaluation face two major limitations. First, creating high-quality project-level evaluation datasets requires extensive domain expertise, leading to prohibitive annotation costs and limited diversity. Second, while recent Agent-as-a-Judge paradigms address the rigidity of traditional unit tests by enabling flexible metrics, their reliance on In-Context Learning (ICL) with general LLMs often results in inaccurate assessments that misalign with human standards. To address these challenges, we propose an agent-driven benchmark construction pipeline that leverages human supervision to efficiently generate diverse project-level tasks. Based on this, we introduce PRDBench, comprising 50 real-world Python projects across 20 domains, each with structured Product Requirement Documents (PRDs) and comprehensive criteria. Furthermore, to overcome the inaccuracy of general LLM judges, we propose a highly reliable evaluation framework powered by a specialized, fine-tuned model. Based on Qwen3-Coder-30B, our dedicated PRDJudge achieves over 90% human alignment in fixed-interface scenarios. Extensive experiments demonstrate that our suite provides a scalable, robust, and highly accurate framework for assessing state-of-the-art code agents.
@@
Code Agent, Evaluation, Agent-as-a-Judge
doi:
HJFB4234
@@ 1. Introduction
In recent years, code agents have made significant progress, capable of solving increasingly complex programming tasks. From initially focusing on single-file code generation to now advancing towards complete project-level software development, Code Agents are demonstrating powerful capabilities in building comprehensive systems. LLM-based Code Agents such as CursorAgent anysphere2025cursoragent, Claude Code anthropic2025claude37 and Gemini CLI mullen2025geminicli have been widely adopted in real-world development scenarios, accelerating the proliferation of automated programming tools.
As code agents continue to advance, a variety of benchmarks have emerged to evaluate their capabilities. Several works chan2024mlebench; huang2024dacodeagentdatascience; daghighfarsoodeh2025deepbenchdeeplearningbenchmark focus on assessing LLMs and code agents in data science tasks, including subtasks such as data processing and machine learning training. Other benchmarks jimenez2023swe; fu2025corecodebench; vergopoulos2025automatedbenchmarkgenerationrepositorylevel select complex code repositories from GitHub pull requests and transform them, often with minimal human intervention, into testable scenarios. In addition, PaperBench starace2025paperbench evaluates the ability of agents to reproduce paper-level code implementations, providing a more rigorous assessment of their research reproducibility.
@@
SWE-Bench jimenez2023swe Pull Requests 12 Web Crawler 2,294 Unit Test 70.3%
MLEBench chan2024mlebench AI Competition 75 Human 75 Test Set 51.1%
DevAI zhuge2024agent Developement 55 Human 365 Agent 73.0%
PaperBench starace2025paperbench AI Research 20 Human (PhD.) 8,316 Human/LLM 21.0%
PRDBench Engineering Development 50 Agent & Human 1,262 Agent (PRDJudge) 45.5%
Table 1. Comparison between PRDBench and project-level code agent benchmarks. We employ the performance of code agents with Claude as a benchmark indicator for assessing the task difficulty of our proposed evaluation suite. The results from external report anthropic2025claude37 is denoted by *, from our implementation is denoted by $\dagger$, otherwise we use the best results from the original paper with author specified Claude-driven agents.
Despite these advancements, current benchmarks face two major limitations. First, the creation of high-quality evaluation datasets requires substantial human effort and expertise. Generating reliable unit tests or grading tasks often depends on expert annotators, especially when migrating real-world scenarios such as Kaggle competitions or GitHub pull requests. These tasks demand domain-level verification and annotation to produce usable test cases. As code agents advance rapidly, the requirements for annotators’ expertise and time investment continue to increase. For example, PaperBench starace2025paperbench recruited ICML authors as annotators, with each annotation task demanding several days of work from PhD-level experts. This process not only incurs high annotation costs, but also restricts the diversity of evaluation data, as recruiting experts from various domains is challenging and often leads to datasets being sourced from a single domain.
Second, the rigidity of the evaluation limits the broad applicability of current benchmarks. Existing benchmarks jimenez2023swe primarily rely on unit test pass rates to assess code agent performance. While unit tests are effective for verifying specific functions or components, they are limited in scope and overly restrictive, enforcing strict requirements on project interfaces and implementation details. To address this rigidity and introduce broader validation methods, recent works have explored LLM-as-a-Judge or Agent-as-a-Judge paradigms. However, these approaches predominantly rely on In-Context Learning (ICL), using zero-shot or few-shot prompts to guide general-purpose LLMs for specific scoring tasks. Relying solely on ICL for complex, project-level evaluation often leads to inaccurate assessments. General LLMs struggle to correctly interpret intricate execution logs or subtle file differences, resulting in significant misalignment with human judgments. For instance, the LLM judger introduced in PaperBench starace2025paperbench achieves a maximum human alignment rate of only 83%. Thus, achieving both evaluation flexibility and high scoring accuracy remains a critical challenge.
@@
Based on this approach, we build a Product Requirement Document (PRD)-centered benchmark, named PRDBench [^1] Source of PRDBench is available in https://github.com/AGI-Eval-Official/PRDBench.. PRDBench consists of 50 real-world coding tasks, each defined by a structured PRD, a well-specified and verifiable criteria scheme, and a standard solution code repository. The coding tasks are sourced from real-world requirements, academic projects, and thesis work, spanning 20 common subdomains. For each task, the PRD provides a criteria scheme that facilitates comprehensive human-like quality assurance (QA) checks. We present a comparison between PRDBench and previous code agent benchmarks in Table 1, where PRDBench offers a comprehensive, multifaceted, project-level benchmark for code agents. During evaluation, we employ the Agent-as-a-Judge paradigm to score code according to the criteria scheme, enabling the evaluation of a wide range of test types beyond unit tests. However, relying on off-the-shelf general LLMs as judges often leads to instability and hallucinations in complex project-level assessments. To ensure highly reliable evaluation, we further introduce a specialized, fine-tuned PRDJudge based on Qwen3-Coder-30B. By training on a curated dataset of evaluation trajectories, our dedicated PRDJudge overcomes the limitations of general models, achieving an over 90% human alignment rate in fixed-interface scenarios.
@@
  • We adopt the Agent-as-a-Judge paradigm powered by a specialized fine-tuned model. Based on Qwen3-Coder-30B, our PRDJudge achieves over 90% human alignment in fixed-interface scenarios, enabling flexible, human-like QA assessment beyond traditional unit tests.
  • We conduct extensive experiments on PRDBench, detailing the training process and data construction of the evaluation agent, and providing useful insights into both state-of-the-art code agents and the evaluation paradigm.
@@ 2.1. Code Agent Evaluation
To keep pace with the rapid development of LLMs’ and agents’ coding capabilities, recent benchmarks increasingly emphasize the construction of executable and complex tasks within real software projects. These benchmarks are typically built by mining human-generated data from online platforms or through extensive manual annotation. Some benchmarks li2024promptinglargelanguagemodels; huang2024dacodeagentdatascience require annotators to manually create test points for each step, while others jimenez2024swebenchlanguagemodelsresolve; yang2024execrepobenchmultilevelexecutablecode; pan2024codevbenchllmsunderstanddevelopercentric; wu2024repomasterevalevaluatingcodecompletion; hai2025impactscontextsrepositorylevelcode rely on authentic submission records, using the inherent difficulty of submitted code as a basis for evaluation. Additionally, some benchmarks li2024evocodebenchevolvingcodegeneration; zhuo2025bigcodebenchbenchmarkingcodegeneration; peng2024humanevalxlmultilingualcodegeneration leverage existing resources and require expert-driven rewriting, which demands significant manual effort. As code agents become more capable, these annotation paradigms require increasingly specialized expertise, making the recruitment of expert annotators for each domain costly and limiting the scalability and diversity of benchmark construction.
In terms of evaluation metrics, most existing benchmarks adopt relatively narrow criteria, such as unit test pass rates jimenez2023swe; li2024promptinglargelanguagemodels; xiang2025scireplicatebenchbenchmarkingllmsagentdriven, which primarily assess functional correctness. ProjectEval liu2025projectevalbenchmarkprogrammingagents simulates user interaction to evaluate projects, focusing on specific types of tasks. However, these evaluation criteria are limited to specific task types and cannot comprehensively assess diverse, complex software projects.
Figure 2. A task example of PRDBench.
@@ 2.2. Agent-as-a-Judge
Due to the inherent limitations of unit test-based benchmarks, recent studies have begun to explore the use of LLMs as judges for agent evaluation, a paradigm often referred to as LLM-as-a-Judge. LLMs can provide capabilities such as image recognition chen2024mllm, code alignment weyssow2024codeultrafeedbackllmasajudgedatasetaligning, and question-answering ho2025llm, breaking the restriction that benchmark metrics must be executable. With the emergence of agents equipped with tools, LLMs are able to read large volumes of files and perform extensive analyses, offering new possibilities for evaluation. DevAI zhuge2024agent is the first work to apply the agent-as-a-judge paradigm to code agent evaluation. It focuses on checking program execution and output format, but lacks comprehensive assessment of overall engineering effectiveness. Some studieswang2025codevisionaryagentbasedframeworkevaluating, have explored LLM-based agent frameworks for evaluating code generation, improving the accuracy of Agent-as-a-Judge through architectural innovations. However, the lack of well-defined metrics and code agents has limited the use of agent-as-a-judge. PRDBench fills this gap with structured criteria and dedicated tools, enabling more robust evaluation.
2.3. Fine-tuned LLMs as Judges
Recent studies suggest that for most tasks, evaluating outputs is inherently easier than generating them together2024finetuning, motivating a line of work that fine-tunes smaller open-source LLMs as dedicated judges. PandaLM wang2024pandalmautomaticevaluationbenchmark, Prometheus kim2024prometheusinducingfinegrainedevaluation, JudgeLM zhu2025judgelmfinetunedlargelanguage, and Auto-J li2023generativejudgeevaluatingalignment demonstrate that fine-tuned models of 7B–33B parameters can match or even surpass GPT-4 on targeted evaluation tasks, while offering significant advantages in cost, speed, and privacy. Notably, Huang et al. huang2025empiricalstudyllmasajudgellm further reveal that such fine-tuned judges essentially operate as highly effective task-specific classifiers, consistently achieving state-of-the-art performance on in-domain benchmarks. These findings motivate our approach of fine-tuning a dedicated PRDJudge specifically for project-level code evaluation, where domain-specific expertise is critical for accurate and reliable scoring.
@@ 3.1. Seed Tasks
Our seed tasks are carefully curated from three primary sources to ensure diversity and real-world relevance: (1) real-world user development requests collected from an internal AI product development platform; (2) publicly available final projects from Computer Science (CS) courses hosted on GitHub; and (3) code reproduction tasks derived from academic theses in the CS domain. To ensure the suitability and consistency of tasks for PRDBench, we apply a rigorous filtering process. Specifically, we require that each selected task can be fully implemented in Python and that all its associated datasets are publicly accessible. We choose Python as the programming language for PRDBench primarily due to its versatility and comprehensive ecosystem, which supports a wide range of programming paradigms and application domains. This ensures that PRDBench tasks are representative of diverse, real-world scenarios.
@@ 3.2. Agent-Driven Data Production
Step 1: PRD and Test Plan Initialization. After selecting seed tasks, we utilize code agents (such as Claude Code) to generate detailed and standardized PRD documents, ensuring clarity and completeness in task specification. The PRD serves as the evaluation blueprint for PRDBench, and includes sections such as Requirement Overview, Functional Requirements, and Data Requirements. Based on the PRD, we employ GPT-4.1 to generate a corresponding metric outline, structuring test cases using the Arrange-Act-Assert (AAA) methodology wei2025developers. The Arrange step sets up the test case by preparing the necessary files, input data, and environment configurations. The Act step focuses on the core behavior to be tested, such as running the program or performing interaction tests to obtain output results. The Assert step verifies the expected outcomes by checking the system’s response or state, ultimately determining whether the test passes or fails. This structure is applicable to a wide range of code-related testing scenarios, providing PRDBench with a executable testing plan.
Figure 3. Domain distribution of PRDBench.
@@
Step 2: Scaffold and Criteria Generation. After initializing with the PRDs and metric outlines, we employ state-of-the-art code agents to generate the code scaffold for each task. The scaffold includes both module design and interface design, serving as the core framework of the entire project. In some cases, particularly for complex algorithms, the implementation may not be fully correct as long as it does not affect the annotation process. Building upon the scaffold, we further utilize code agents to expand and refine the metric outline into a specific criteria scheme, which includes necessary test interfaces and test artifacts. The presence of well-defined code interfaces in the scaffold simplifies and standardizes the generation of the criteria scheme, making the process more efficient and consistent.
@@
Step 3: Human Inspection. With the scaffold and the criteria scheme in place, the inspection process for human annotators becomes straightforward. Annotators conduct inspection only by running the tests to verify whether the interfaces function correctly and whether the expected outputs in the criteria scheme align with the PRD requirements. Notably, even for tasks sourced from expertise projects or academic papers, annotators only need basic computer science knowledge to perform effective annotation.
@@
Step 4: Agent-Based Fix and Refinement. If any issues are identified during step 3, annotators shall provide targeted feedback to the code agent for refining the scaffold or criteria scheme. The code agent then revises the relevant components, and the inspection process is repeated. This iterative process continues until all issues are resolved. To ensure the rigorous quality and sufficient complexity of our benchmark, we establish a strict inclusion criterion: only tasks that have undergone at least 5 rounds of such iterative human-agent refinement are retained.
@@
Step 5: Remove Scaffold. Finally, the scaffold is removed, retaining only the PRD, criteria scheme, test artifacts, and data. This forces the evaluated agents to generate code from scratch, enabling a rigorous assessment of their end-to-end development capabilities.
@@
3.3. Data Statistics
PRDBench comprises 50 project-level Python tasks that span 20 distinct application domains. To ensure broad and balanced coverage, we utilized GPT-4 to assign domain labels to all candidate tasks before strategically sampling the final dataset. As illustrated in Figure 3, we include at least one representative task from less common domains. For more prevalent areas, such as data processing and machine learning, we select tasks that demand a diverse array of technical approaches and sub-skills. This sampling strategy ensures that PRDBench accurately represents diverse, real-world Python development scenarios.
Figure 4. Scaffold line counts distribution.
These tasks exhibit substantial complexity. On average, each structured PRD contains 105.22 lines of detailed requirements. Furthermore, Figure 4 presents the distribution of scaffold line counts recorded during the annotation process, highlighting the realistic scale of the projects.
For comprehensive evaluation, PRDBench encompasses a total of 1,258 evaluation metrics. These metrics collectively simulate the different stages of quality assurance (QA) in real-world software engineering and are categorized into three main types: